you and a friend go to a mexican restaurant. you order two tacos and three enchiladas, and your friend orders three tacos and five enchiladas. your bills is $7.80 plus tax, and your friends bill is $12.70 plus tax. how much does each taco and each enchilada cost?
Answer by jorel1380(2518)

2t+3e=7.80
3t+5e=12.70
--------------
6t+9e=23.40
6t+10e=25.40
e=2.00
t=.90
Tacos are $.90 each. Enchiladas are $2.00 each. Don't forget to tip..